# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v2

@@ VI.1.2 Mixed ESS
The above arguments using Lyapunov function, $V_{1}$, carry over to the mixed ESS’s under and- and sum-definitions as well, leading to the conclusion: Mixed ESS (except mixed 2ESS) is always locally asymptotically fixed point of replicator equation. The case of 2ESS (or-definition) is interesting.
First, we recall that interior fixed point can only be a center or saddle node in bimatrix games (see Sec. III.3). We now show that— extending the argument presented in Ref. Hofbauer and Sigmund (1998) for two player setting—similar assertion holds true even for multiplayer bihypermatrix games, i.e., an interior fixed point can neither be a source or sink. It is straightforward to see that, by definition, the internal fixed point ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) of Eq. (71) with ${\sf A}={\sf D}=0$ satisfies the following condition:
@@
On linearizing the replicator equation about the fixed point, takes the form
@@
Similarly, we can prove that ${\partial\dot{y}_{j}}/{\partial y_{k}}=0$. Now, the characteristic polynomial, $p(\lambda)$ (say), can be obtained as $\det({\sf J}-\lambda I)$. It is evident that $p(\lambda)=(-1)^{(m-1)+(n-1)}p(-\lambda)$—a result derived by changing the signs of the first $m-1$ columns and the last $n-1$ rows of $\det({\sf J}-\lambda I)$. Hence, if $\lambda$ is an eigenvalue of ${\sf J}$, then $-\lambda$ must also an eigenvalue, thereby preventing the interior fixed point from being sink or source.
Next, we come to mixed 2ESS in asymmetric hypermatrix game. One could expect that as done previously in Sec. IV.2, $V_{r}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ as a candidate for Lyapunov function. For this function to be a valid Lyapunov function one needs an analog of Theorem 2 for multiplayer games, that is, the following proposition needs to be true:
@@
$\mu{\sf A}[\mathbf{\hat{x}};\mathbf{x},\mathbf{x}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y},\mathbf{y}]-\mu{\sf A}[\mathbf{x};\mathbf{x},\mathbf{x}]-{\sf B}\left[\mathbf{x};\mathbf{y},\mathbf{y}\right]$ $=$ $-5\delta_{x}\delta_{y}+\mathcal{O}(\delta_{x}\delta_{y}\mu),$ (86a)
${\sf C}[\mathbf{\hat{y}};\mathbf{x},\mathbf{x}]+\mu{\sf D}[\mathbf{\hat{y}};\mathbf{y},\mathbf{y}]-{\sf C}\left[\mathbf{y};\mathbf{x},\mathbf{x}\right]-\mu{\sf D}[\mathbf{y};\mathbf{y},\mathbf{y}]$ $=$ $+7\delta_{y}\delta_{x}+\mathcal{O}(\delta_{y}\delta_{x}\mu).$ (86b)
Since signs of $\delta_{x}$ and $\delta_{y}$ can be positive or negative, at least one of the R.H.S.’s of Eq. (86) is positive in the entire neighbourhood; thus, the criteria of or-definition is satisfied but it does not meet the criteria of ESS of and- and sum-definitions. Thus, the locally asymptotically stable fixed point given by Eq. (84) is 2ESS. But does it satisfy inequality (81)? As we see below, it does not (when $\mu$ is very small as assumed in the calculations above).
@@ VII Conclusions
The primary objective of this study has been to critically revisit various definitions of ESS, specifically in the context of asymmetric (involving both inter and intra-specific interactions) and bimatrix games (with solely inter-specific interactions), and to unravel their suitability. Through our analyses, we have brought forward intricacies inherent in the definitions and have suggested their corresponding resolutions by examining them through the lens of dynamical stability. The insights, thus, gained resolve the prevailing ambiguities surrounding the ESS concept on asymmetric interaction scenarios.
@@
As we have described in this paper, the definitions of ESS in a population that involves two subpopulations can be broadly classified into three classes, which we may appositely call (i) and-definitions, (ii) sum-definitions, and (iii) or-definitions. In the case of and-definitions (Definitions 1a to 1e), the fitness of host individuals in each subpopulation is compared against that of the corresponding mutant. Whereas, in the case of sum-definitions (Definitions 2a to 2e) one compares the added fitnesses of residents in both the subpopulations with the combined fitness of corresponding mutants. Finally, in the case of or-definitions (Definitions 3a to 3e), a state is said to be ESS if any one subpopulation’s residents are fitter against the mutants. The ESS under or-definition has a special name in the literature: 2ESS Cressman and Tao (2014) which, in the case of bimatrix games, is equivalent to another concept called Nash–Pareto pair Hofbauer and Sigmund (1998).
@@
If no intra-specific interaction is present, one has so called bimatrix or bihypermatrix games. One finds that ESS’s of and- and sum-definitions can only be pure, whereas 2ESS (ESS of or-definitions) can be either pure or mixed. However, when it comes to the desirable property that an ESS better be locally asymptotically stable fixed point of the appropriate replicator equation, it may not possess the property if it is non-strict NE. When intra-specific interactions are present in the so-called asymmetric matrix or hypermatrix games, ESS’s under all three classes of definitions can be either pure or mixed. While it seems that mixed ESS’s are locally asymptotically stable fixed points of the corresponding replicator equation, pure ESS need not be so if it corresponds to non-isolated fixed points. Also, we bring again to the readers’ attention that apart from using information-theoretic idea in finding the Lyapunov function, we have used the Fisher information metric in the form of the Shahshahani metric to decipher an interpretation of the weight factor ($r$) present in an equivalent definition of 2ESS: $r$ is the inverse of the rescaling factor in a distinct class of games known as rescaled partnership games.
@@
The most important take-home message, we would like to convey is that while strict NE is ESS under all three classes of definitions, other ESS’s need not simultaneously satisfy all the classes. Not only that, the latter ESS’s may not always be locally asymptotically stable under replicator equation. In a given situation, it may be practical to adopt that definition which renders ESS locally asymptotically stable; after all, if an ESS is not reached dynamically starting from a mutant-invaded state, then it doesn’t seem to serve its purpose. The asymptotic approach towards ESS state is also compatible with the principle of minimum relative entropy—a rather general principle of nature.
@@
We conclude by highlighting potential avenues for extending the findings of this paper. While we have provided an illustrative example, the rigorous proof of results concerning mixed 2ESS in asymmetric hypermatrix games remains an open question. Establishing pure 2ESS’s asymptotic stability in multiplayer games using techniques beyond linear stability analysis is also missing. Moreover, since the study in this paper has been confined to matrix games, the implications of nonlinear payoffs Archetti and Scheuring (2012); Patra et al. (2022) in non-matrix population games remains unexplored. Exploring the realm of asymmetric games in finite Nowak et al. (2004); Sekiguchi and Ohtsuki (2015) populations to locate an appropriate definition of ESS is also an interesting problem. Finally, extending the concept of ESS to non-convergent outcomes, which may manifest as oscillatory orbits rather than fixed points Bhattacharjee et al. (2023), opens up another avenue for investigation.
@@ Acknowledgements.
Appendix A Comment on Ref. Cressman (1996)’s proof of Theorem 1
In Appendix A of Ref. Cressman (1996), Theorem 1 is seemingly proven in the context of asymmetric games which are defined to include bimatrix games. While the elaborate proof is best seen in the original reference, let us briefly point out at which point of argument the proof ceases to be valid for bimatrix games.
@@
We begin by referring to a statement present in the proof: “When $\theta=0$, the point is$-((p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1}),(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$ which is in quadrant IV. Similarly, when $\theta=\pi/2$ the point is in quadrant II.” Here $(p^{*},q^{*})$ is the resident state and $(p_{1},q_{1})$ is the mutant state, which in our notation denoted as $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ and $(\mathbf{x},\mathbf{y})$, respectively. $A,\,B,\,C,\,D$ are payoff matrices ${\sf A,\,B,\,C,\,D}$. In this proof, one tries to show that Definition 3d implies inequality (46); for this purpose, one introduces $\theta$ as a parameter to explore in $\mathbb{R}^{2}$ space spanned by $(p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1})$ and $(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$ as the two coordinates. Idea is to basically find a point in this $\mathbb{R}^{2}$ space at which Definition 3d is not satisfied and hence, to contradict an initial hypothesis that there is no such $r$ for which inequality (46) is satisfied.
@@
A subsequent sentence in the scheme of proof is: “Thus, for some $\theta_{0}$ between 00 and $\pi/2$, the point has zero first component and, by Definition 4.3, negative second. Then for $\theta_{0}+\pi/2$, the point has zero first component and positive second. This gives our contradiction.” Here Definition 4.3 is Definition 3d of this paper. For bimatrix games, however, $A=0$ and $D=0$, and since this proof is valid for interior fixed point (which is also a mixed NE), we can add $(p^{*}-p_{1})\cdot Bq^{*}$ (which is zero) to the coordinate, $(p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1})$, and $(q^{*}-q_{1})\cdot Cp^{*}$ (which also is zero) to the second coordinate, $(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$. Subsequently, on redefining $(p^{*}-p_{1})$ as $\bm{\xi}$ and $(q^{*}-q_{1})$ as $\bm{\eta}$, the corresponding point in $\mathbb{R}^{2}$ takes the following form: $(\bm{\xi}\cdot B\bm{\eta},\bm{\eta}\cdot C\bm{\xi})$. To complete the proof, one needs to argue that the point has zero first component and negative second component—this is exactly what is not valid because if first component is zero then automatically second component is also zero. Let us see explicitly why this is so.
@@
We know that any $2\times 2$ bimatrix game can be expressed in the form of either a rescaled zero-sum game or a rescaled partnership game Hofbauer and Sigmund (1998). For the rescaled zero-sum game, ${\sf C}=-c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$; whereas for the rescaled partnership game, we have ${\sf C}=c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$ (see Section V). For a rescaled zero-sum game or rescaled partnership game, easy calculations show that if $\bm{\xi}\cdot B\bm{\eta}=0$ then $\bm{\eta}\cdot C\bm{\xi}=0$ Hofbauer and Sigmund (1998). This is because $\sum_{i}\xi_{i}=\sum_{i}\eta_{i}=0$ and hence, we can trivially show: $\bm{\xi}\cdot\Omega\bm{\eta}=\bm{\eta}\cdot\Omega^{\prime}\bm{\xi}=0$.
In conclusion, if one accepts the theorem—present in Ref. Cressman (1996)—exactly as worded, then one is prone to erroneously deducing that mixed 2ESS in bimatrix games is asymptotically stable. This oversight in Ref. Cressman (1996) extends beyond just the afore-discussed $2\times 2$ case, simply because it is known that asymptotic stability of an interior fixed point (of corresponding replicator equation) is not possible for bimatrix games in general Hofbauer and Sigmund (1998).
Appendix B ESS under Definition 2e for bihypermatrix games can only be pure

### v3

@@ VI.1.2 Mixed ESS
The above arguments using Lyapunov function, $V_{1}$, carry over to the mixed ESSes under and- and sum-definitions as well, leading to the conclusion: Mixed ESS (except mixed 2ESS) is always locally asymptotically fixed point of the replicator equation. The case of 2ESS (or-definition) is interesting.
First, we recall that interior fixed point can only be a center or saddle node in bimatrix games (see Sec. III.3). We now show that—extending the argument presented in Ref. Hofbauer and Sigmund (1998); Schuster et al. (1981) for two player setting—a similar assertion holds true even for multiplayer bihypermatrix games, i.e., an interior fixed point can neither be a source or sink. It is straightforward to see that, by definition, the internal fixed point ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) of Eq. (68) with ${\sf A}={\sf D}=0$ satisfies the following condition:
@@
On linearizing the replicator equation about the fixed point, yields the Jacobian matrix
@@
Similarly, we can prove that ${\partial\dot{y}_{j}}/{\partial y_{k}}=0$. Now, the characteristic polynomial, $p(\lambda)$ (say), can be obtained as $\det({\sf J}-\lambda I)$. It is evident that $p(\lambda)=(-1)^{(m-1)+(n-1)}p(-\lambda)$—a result derived by changing the signs of the first $m-1$ columns and the last $n-1$ rows of $\det({\sf J}-\lambda I)$. Hence, if $\lambda$ is an eigenvalue of ${\sf J}$, then $-\lambda$ must also be an eigenvalue, thereby preventing the interior fixed point from being sink or source.
Next, we come to mixed 2ESS in asymmetric hypermatrix games. One could expect that as done previously in Sec. IV.2, $V_{r}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ as a candidate for Lyapunov function. For this function to be a valid Lyapunov function one needs an analog of Theorem 1 for multiplayer games, that is, the following proposition needs to be true:
@@
[…]
Since signs of $\delta_{x}$ and $\delta_{y}$ can be positive or negative, at least one of the R.H.S.’s of Eq. (83) is positive in the entire neighbourhood; thus, the criteria of or-definition is satisfied but it does not meet the criteria of ESS of and- and sum-definitions. Thus, the locally asymptotically stable fixed point given by Eq. (81) is a 2ESS. But does it satisfy inequality (78)? As we see below, it does not (when $\mu$ is very small as assumed in the calculations above).
@@ VII Conclusions
The primary objective of this study has been to critically revisit various definitions of ESS, specifically in the context of asymmetric (involving both inter and intra-specific interactions) and bimatrix games (with solely inter-specific interactions), and to unravel their suitability. Through our analyses, we have brought forward intricacies inherent in the definitions and have suggested their corresponding resolutions by examining them through the lens of dynamical stability. The insights thus gained help resolve the prevailing ambiguities surrounding the ESS concept in asymmetric interaction scenarios.
@@
As we have described in this paper, the definitions of ESS in a population that involves two subpopulations can be broadly classified into three classes, which we may appositely call (i) and-definitions, (ii) sum-definitions, and (iii) or-definitions. In the case of and-definitions (Definitions 1a to 1e), the fitness of host individuals in each subpopulation is compared against that of the corresponding mutant. Whereas, in the case of sum-definitions (Definitions 2a to 2e) one compares the added fitnesses of residents in both the subpopulations with the combined fitness of corresponding mutants. Finally, in the case of or-definitions (Definitions 3a to 3e), a state is said to be ESS if at least one of the two subpopulations of residents are fitter against the mutants. The ESS under or-definition has a special name in the literature: 2ESS Cressman and Tao (2014).
@@
If no intra-specific interaction is present, one has so called bimatrix or bihypermatrix games. One finds that ESSes of and- and sum-definitions can only be pure, whereas 2ESS (ESS of or-definitions) can be either pure or mixed. However, when it comes to the desirable property that an ESS should be locally asymptotically stable fixed point of the appropriate replicator equation, it may not possess the property if it is a non-strict NE. When intra-specific interactions are present in the so-called asymmetric matrix or hypermatrix games, ESSes under all three classes of definitions can be either pure or mixed. While it seems that mixed ESSes are locally asymptotically stable fixed points of the corresponding replicator equation, pure ESS need not be so if it corresponds to non-isolated fixed points. Also, we bring again to the readers’ attention that apart from using information-theoretic idea in finding the Lyapunov function, we have used the Fisher information metric in the form of the Shahshahani metric to decipher an interpretation of the weight factor ($r$) present in an equivalent definition of 2ESS: $r$ is the inverse of the rescaling factor in a distinct class of games known as rescaled partnership games.
@@
The most important take-home message, we would like to convey is that while a strict NE is an ESS under all three classes of definitions, other ESSes need not simultaneously satisfy all the classes. Not only that, the latter ESSes may not always be locally asymptotically stable under replicator equation. In a given situation, it may be practical to adopt that definition which renders ESS locally asymptotically stable; after all, if an ESS is not reached dynamically starting from a mutant-invaded state, then it doesn’t seem to serve its purpose. The asymptotic approach towards ESS state is also compatible with the principle of minimum relative entropy—a rather general principle of nature.
@@
We conclude by highlighting potential avenues for extending the findings of this paper. While we have provided an illustrative example, the rigorous proof of results concerning mixed 2ESS in asymmetric hypermatrix games remains an open question. Establishing the asymptotic stability of a pure 2ESS in multiplayer games using techniques beyond linear stability analysis is also missing. Moreover, since the study in this paper has been confined to matrix games, the implications of nonlinear payoffs Archetti and Scheuring (2012); Patra et al. (2022) in non-matrix population games remains unexplored. Exploring the realm of asymmetric games in finite Nowak et al. (2004); Sekiguchi and Ohtsuki (2015) populations to locate an appropriate definition of ESS is also an interesting problem. Finally, extending the concept of ESS to non-convergent outcomes, which may manifest as oscillatory orbits rather than fixed points Bhattacharjee et al. (2023); Dubey et al. (2024), opens up another avenue for investigation.
@@ Acknowledgements.
Appendix A Alternative definitions of ESS in bimatrix games
In this section, we present alternative ESS definitions that, although not equivalent to those in the main text, are motivated by the same underlying intent.
A.1 Alternative and-definitions
We write three equivalent alternative and-definitions in the context of bimatrix games as follows:
  1. Definition 1a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\in\Sigma_{m}\times\Sigma_{n}$ there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$, with $0<\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}}),\leavevmode\nobreak\ \mathbf{{x}}\neq\mathbf{\hat{x}}\leavevmode\nobreak\ \text{\emph{and}}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}),\leavevmode\nobreak\ \mathbf{{y}}\neq\mathbf{\hat{y}}.$ (84)
  2. Definition 1b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})$ the following conditions hold,
[…]
    (ii) and if $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, then for all $\mathbf{y}$,
    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y};$ (86)
    (iii) and if $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$, then, for all $\mathbf{x}$,
    $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (87)
    As written, conditions (ii) and (iii) are extremely restrictive: They can never be satisfied. This is because, if $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, then a mutant of the form $(\mathbf{x},\mathbf{\hat{y}})$ renders inequality (86) invalid. Therefore, the above definition effectively reduces to the following stricter and simpler form, as given in Hofbauer and Sigmund (1988).
  3. Definition 1b′′: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})$ the following conditions hold,
[…]
    Finally, we write the neighborhood version of this definition:
  4. Definition 1c: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that
    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y},\leavevmode\nobreak\ \mathbf{{x}}\neq\mathbf{\hat{x}}\leavevmode\nobreak\ \text{\emph{and}}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x},\leavevmode\nobreak\ \mathbf{{y}}\neq\mathbf{\hat{y}}.$ (89)
A.2 Nash-Pareto pair
Hofbauer and Sigmund proposed an alternative definition of ESS to accommodate mixed ESS in bimatrix games Hofbauer and Sigmund (1998, 2003). They refer to this concept as the Nash–Pareto pair:
  1. Definition 3b: A state $(\hat{\mathbf{x}},\hat{\mathbf{y}})$ is a Nash–Pareto pair if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold:
[…]
    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which both inequalities in (i) hold as equalities, following condition holds:
    $\text{if}\leavevmode\nobreak\ \mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}<\mathbf{x}\cdot{\sf B}\mathbf{y}\leavevmode\nobreak\ \text{then}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\text{ and if}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}<\mathbf{y}\cdot{\sf C}\mathbf{x}\leavevmode\nobreak\ \text{then}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}.\$ (91)
We would like to bring it to the fore that the Nash–Pareto pair is different from 2ESS (equivalently, Definition 3b). Observe that, we can rewrite (91) as ‘either $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}\geq\mathbf{x}\cdot{\sf B}\mathbf{y}$ or $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ and either $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}\geq\mathbf{y}\cdot{\sf C}\mathbf{x}$ or $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$’. This follows from the material implication rule Hurley (2014): ‘if $P$ then $Q$’ is equivalent to ‘either not P or Q’ (i.e., symbolically, $P\rightarrow Q\equiv\neg P\lor Q$). Thus, even if the resident state satisfies inequalities in (90) as equalities and furthermore, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}=\mathbf{x}\cdot{\sf B}\mathbf{y}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}=\mathbf{y}\cdot{\sf C}\mathbf{x}$ hold true, then $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ still qualifies as a Nash-Pareto pair; however, clearly, such an $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is not an ESS under or-definition.
Appendix B Non-applicability of Lyapunov function (44) in bimatrix games
The Lyapunov function given in Eq. (44) appears quite general, and one might think that it applies to bimatrix games as well. However, as we already know that the internal fixed point in bimatrix games cannot be asymptotically stable under replicator dynamics. Therefore, the Lyapunov function in Eq. (44) is not applicable in the case of bimatrix games. In this section, we discuss why this is the case. To do so, we use the proof of Theorem 1 provided in Appendix A of Ref. Cressman (1996). While the elaborate proof is best seen in the original reference, let us briefly point out at which point of the argument in the proof ceases to be valid for bimatrix games.
@@ Appendix B Non-applicability of Lyapunov function (44) in bimatrix games
We begin by referring to a statement present in the proof: “When $\theta=0$, the point is$-((p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1}),(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$ which is in quadrant IV. Similarly, when $\theta=\pi/2$ the point is in quadrant II.” Here $(p^{*},q^{*})$ is the resident state and $(p_{1},q_{1})$ is the mutant state, which in our notation denoted by $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ and $(\mathbf{x},\mathbf{y})$, respectively. $A,\,B,\,C,\,D$ are payoff matrices ${\sf A,\,B,\,C,\,D}$. In this proof, one tries to show that Definition 3d implies inequality (43); for this purpose, one introduces $\theta$ as a parameter to explore in $\mathbb{R}^{2}$ space spanned by $(p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1})$ and $(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$ as the two coordinates. The idea is to basically find a point in this $\mathbb{R}^{2}$ space at which Definition 3d is not satisfied and hence, to contradict an initial hypothesis that there is no such $r$ for which inequality (43) is satisfied.
@@
A subsequent sentence in the scheme of the proof is: “Thus, for some $\theta_{0}$ between 00 and $\pi/2$, the point has zero first component and, by Definition 4.3, negative second. Then for $\theta_{0}+\pi/2$, the point has zero first component and positive second. This gives our contradiction.” Here Definition 4.3 is Definition 3d of this paper. For bimatrix games, however, $A=0$ and $D=0$, and since this proof is valid for interior fixed point (which is also a mixed NE), we can add $(p^{*}-p_{1})\cdot Bq^{*}$ (which is zero) to the coordinate, $(p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1})$, and $(q^{*}-q_{1})\cdot Cp^{*}$ (which also is zero) to the second coordinate, $(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$. Subsequently, on redefining $(p^{*}-p_{1})$ as $\bm{\xi}$ and $(q^{*}-q_{1})$ as $\bm{\eta}$, the corresponding point in $\mathbb{R}^{2}$ takes the following form: $(\bm{\xi}\cdot B\bm{\eta},\bm{\eta}\cdot C\bm{\xi})$. To complete the proof, one needs to argue that the point has zero first component and negative second component—this is exactly what is not valid because if the first component is zero then the second component is automatically zero as well. Let us see explicitly why this is so.
@@
We know that any $2\times 2$ bimatrix game can be expressed in the form of either a rescaled zero-sum game or a rescaled partnership game Hofbauer and Sigmund (1998). For the rescaled zero-sum game, ${\sf C}=-c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$; whereas for the rescaled partnership game, we have ${\sf C}=c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$ (see Section V). For a rescaled zero-sum game or a rescaled partnership game, easy calculations show that if $\bm{\xi}\cdot B\bm{\eta}=0$ then $\bm{\eta}\cdot C\bm{\xi}=0$ Hofbauer and Sigmund (1998). This is because $\sum_{i}\xi_{i}=\sum_{i}\eta_{i}=0$ and hence, we can trivially show: $\bm{\xi}\cdot\Omega\bm{\eta}=\bm{\eta}\cdot\Omega^{\prime}\bm{\xi}=0$.
In conclusion, due to the above arguments, the equivalence between inequality (43) and 2ESS does not hold in bimatrix games, which prevents Eq. (44) from being a valid Lyapunov function in such games. Finally, we remark that the above arguments extends beyond $2\times 2$ case, simply because it is known that asymptotic stability of an interior fixed point (of corresponding replicator equation) is not possible for bimatrix games in general Hofbauer and Sigmund (1998).
Appendix C ESS under Definition 2e for bihypermatrix games can only be pure